Crew 260 Operations Report November 1st

MDRS Operations Report 1Nov2021

Name of person filing report: Atila Meszaros

Reason for Report: Daily

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— SOC %: 74 at 19:00

Notes on solar: Nothing to report

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 30 gallons

Water Meter: 0152873.8

Water (static tank) – 350 gallons

Static to Loft Pump used – yes

Water in GreenHab – 90 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 90 gallons

Hab toilet tank emptied: No

Perseverance rover used: Yes

Hours: 220.6

Beginning charge: 100

Ending charge: 39

Currently charging: No

Sojourner rover used: No

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Spirit rover used: Yes

Hours: 141.0

Beginning charge: 100

Ending charge: 75

Currently charging: No

Opportunity rover used: No

Hours: 73.0

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Curiosity rover used: No

Hours: 148.0

Beginning charge: 100

Ending charge: 75

Currently charging: Yes

Notes on rovers: We cleaned the battery contacts with baking soda and water and refilled the cells with 5-6 gallons of destilled water. We need to find a cord for Percy cause it’s melting.

ATV’s Used: (Honda, 350.1, 350.2, 350.3): Ran 350’s 1 and 2

Reason for use: Training

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: for 30 minutes

Notes on ATVs: Honda still on the jack awaiting tire repair/replacement

General notes and comments:

Summary of internet: All accounts are nominal.

Summary of suits and radios: Suits and radio training for Crew 260. Everyone chose and was asigned their prefered spacesuit. All the one-piece and two-piece suits are currently charging for tomorrow’s first EVA.

Summary of general operations: Nothing to report

Summary of Hab operations: Space Rebound Mission started today. The rest of the team arrived in the morning and we started the program at 09:00. The team received a tour of the station and went trough radios, suit and rover training. We are currently conserving power because we plugged-in the rovers too late, cause the generator is still not online.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: We removed the shade cloth.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.